Gene's car traveled 210 miles averaging a certain speed. If the car had gone 5 mph faster, the trip would have taken 1 hour less. Find the average speed.
Answer by mananth(16949)

speed x
speed x - 5

Distance = same 210 miles
original time – time with increased speed =1
t=d/t
210/x-210/(x+5) =1
LCD= x (x+5)
multiply by LCD
210(x+5)-210x=1x(x+5)
210x+1050-210x=1x^2+5x
1050=1x^2+5x
x^2+5x-1050=0

a= 1 b= 5 c= -1050
x1=(-5+sqrt(25-( -4200) ) / 2
= 4225
x1= 30

x2=(5-sqrt(25-(-4200))/ 2

x2= -35
Ignore negative value
speed of car = 30 mph